OpenStudy (anonymous):

Use the table and number line to answer Parts A and B. Number of Times on Favorite Roller Coaster 12 7 4 20 12 6 13 6 13 6 7 5 11 14 A. Find the minimum, Q1 , Q2, Q3, and the maximum of the data. B. Use the number line to draw a box-and-whisker plot of the data. 0 5 10 15 20 25

OpenStudy (anonymous):

put your number in order.... 4,5,6,6,6,7,7,11,12,12,13,13,14,20 Q2, or the median, is (7 + 11)/2 = 18/2 = 9 Q1 = 6 Q3 = 13 Minimum = 4 Maximum = 20